HETAV PATEL

Grade 11,

Volume of a gas at STP is 1.12×10^–7 cc... Calculate the molecules in it.

Volume of a gas at STP is 1.12×10^–7 cc... Calculate the molecules in it.

Grade:11

1 Answers

Amit kumar
6 years ago
Molar valoume at STP=22400ml.so number of mole=1.12×10^-7/22400=5×10^-12. Number of molecules=5×10^-12×6.022×10^23=3.01×10^12